Suppose you want to buy a tablet that costs P 9,000.00

a. How long would it take you to raise the money if you could save:

a.1. Ten pesos a week?

a.2 Twenty pesos a week?

a.3 Fifty pesos a week?

a.4 One hundred pesos a week?

a.5 Two hundred pesos a week?


c. Write an equation that best represent the situation.

d. Explain what will happen as the weekly savings approaches zero.

d.1 What do you think is the possible lowest weekly savings?

d.2 What do you think is the possible highest weekly savings?


1
Expert's answer
2021-11-08T12:05:47-0500

a.1 "w=\frac{9000}{10}=900\;weeks."

a.2 "w=\frac{9000}{20}=450\;weeks."

a.3 "w=\frac{9000}{50}=180\;weeks."

a.4 "w=\frac{9000}{100}=90\;weeks."

a.5 "w=\frac{9000}{200}=45\;weeks."


c. "w=\frac{9000}{s}"


d. As the weekly savings approaches zero, the number of weeks approaches to infinity.

d.1 The possible lowest weekly savings is P0.01.

d.2 The possible highest weekly savings is P9000.00.